# Explainable Metrics for the Assessment of Neurodegenerative Diseases through Handwriting Analysis

### v2

@@ Abstract
Motor dysfunction is a common sign of neurodegenerative diseases (NDs) such as Parkinson’s disease (PD) and Alzheimer’s disease (AD), but may be difficult to detect, especially in the early stages. In this work, we examine the behavior of a wide array of explainable metrics extracted from the handwriting signals of 113 subjects performing multiple tasks on a digital tablet, as part of the Neurological Signals dataset. The aim is to measure their effectiveness in characterizing NDs, including AD and PD. To this end, task-agnostic and task-specific metrics are extracted from 14 distinct tasks. Subsequently, through statistical analysis and a series of classification experiments, we investigate which metrics provide greater discriminative power between NDs and healthy controls and amongst different NDs. Preliminary results indicate that the tasks at hand can all be effectively leveraged to distinguish between the considered set of NDs, specifically by measuring the stability, the speed of writing, the time spent not writing, and the pressure variations between groups from our handcrafted explainable metrics, which shows $p$-values lower than 0.0001 for multiple tasks. Using various binary classification algorithms on the computed metrics, we obtain up to 87% accuracy for the discrimination between AD and healthy controls (CTL), and up to 69% for the discrimination between PD and CTL.
@@
Handwriting Analysis, Alzheimer’s Disease, Parkinson’s Disease, Digital Biomarkers, Neurological Signals
@@ I Introduction
Neurodegenerative diseases (NDs), such as Parkinson’s disease (PD) and Alzheimer’s disease (AD), have detrimental effects on shared and disparate regions of the brain, resulting in a gradual deterioration of cognitive, motor, and behavioral function. Numerous attempts have been made to track dysfunction utilizing non-invasive, wearable or digital sensors. Handwriting, in particular, with its complex syngergism between cognitive, kinetic, and perceptual motor components, is therefore an obvious avenue to evaluate NDs [1, 2]. Successful handwriting utilizes cortical and subcortical brain regions pertaining to learning, cognitive control, praxis, vigilance and coordination across the cerebral cortex, basal ganglia, and cerebellum [3]. Alterations in its patterns can serve as a source of prominent digital markers, indicative of both the presence and severity of the disease [4], as exemplified by micrographia being a common early sign of PD [5].
Paper and digital handwriting tests [4, 5, 1] have been used to assess the presence of NDs through the dynamic and spacial characteristics of handwriting across sets of tasks. The bulk of prior work studied PD as micrographia was an obvious target [6, 7, 8, 9]. Subjects with PD often suffer from micrographia [10, 11], tremor [11], and bradykinesia [11], a collection of signs collectively referred to as PD dysgraphia[1]. More recently, numerous reports have begun to tackle AD detection through handwriting test, as the condition has grown more common [12].
The handwriting tests used for assessing NDs can be categorized into several subtypes, dependent upon complexity and specific task instructions. The first category consists of simple drawing shapes, such as spirals and straight lines. These simple drawings readily evaluate trajectory, tremor, dimensionality, velocity, or acceleration [13]. The second category involves writing words or short sentences. When conducted on a digital tablet, these tasks permit analysis of in-air movements and motor planning between words [13]. The collection of in-air movements is crucial, as the time spent in-air reflects a person’s preparation and planning for subsequent movements, implicitly indicative of their cognitive processesing [5]. Additionally, in-air and on-surface movements are statistically independent and provide distinct information about the writer [14]. A prolonged in-air time may be indicative of hesitations or impaired motor planning and sequencing. In contrast, fluid handwriting execution with minimal in-air time, demonstrates the presence of an established motor plan. The third category encompasses complex tasks, including drawing complex figures or transcribing paragraphs. The clock drawing task is an established modality [15] to detect cortical and subcortical dysfunction across several NDs that reflect organization and visuospatial skills; or visuoconstruction deficits [16]. Copying tasks provide comparisons of writing variations with respect to different stimuli [17]. The use of a variety of these tasks provides a full profile of a person’s handwriting capabilities.
In this study, we collected a new data set from 103 participants with several NDs: AD, Mild Cognitive Impairment (MCI), PD and Related Disorders (the latter term inclusive of atypical PD and secondary Parkinsonism in which symptoms were initially ascribed to PD but later concluded to be an alternative condition, which we will term [PDM]). This dataset comprises recordings of handwriting on a digital tablet across 14 specific tasks specifically designed to assess all aforementioned NDs. We extract explainable and quantitative endpoints, also called metrics, from 14 handwriting tasks and pinpoint key metrics that distinguish between healthy controls and the NDs subgroups, including an active control group to compare against PD (i.e., PDM). We first explore a more general approach to obtain task-agnostic metrics, independently of their original purpose or content.
Secondly, a task-dependent approach is adopted to fully explore which traits of the participants can be used for discrimination, following the expected behavior that each group of participants is supposed to exhibit for each group of tasks.
Furthermore, we use statistical analysis and machine learning classifiers to compare the prediction power of the extracted metrics using binary comparisons between groups. The results obtained from that analysis highlight the possible value of these endpoints as digital biomarkers.
The main contributions of this article are:
  • The introduction of a handwriting dataset for the assessment of NDs
  • The inclusion of several ND groups, including an active control group to compare to PD (i.e, PDM group), where all precedent studies were more limited in scope.
  • The use of Handwriting for AD and MCI assessment was previously primarily reserved for PD assessment.
  • A large set of task-agnostic and task-specific metrics that show significant differences between groups of participants following expected behaviors.
  • The introduction to inter-tasks comparisons of metrics, to study the behavior differences for a given participants facing a difference in their goals, such as changing the drawing hand.
@@ II-A Data Collection
The authors of this study collected a dataset from 113 participants from a larger digital biomarker cohort, NeuroLogical Signals, which our group has previously reported [18, 19, 20, 21, 22, 23]. These participants were either categorized as healthy (CTL) or diagnosed and treated for a ND by a subspecialist neurologist or geriatric psychiatrist from Johns Hopkins Hospital. Participants diagnosed with NDs were categorized into aforementioned groups based on disease specific clinical diagnostic criteria diagnosis:
@@
  • PD participants met international recognized diagnostic criteria at a ”clinically established” [27] level of certainty and were followed by a movement disorders neurologist.
  • PDM participants presented with one or more core symptoms of PD, meeting a ”Probable PD” diagnosis but which was later revised based on evolving characteristics or supportive testing. This is a mixed, active comparator group includes patients later diagnosed with Dementia with Lewy Bodies, Fahr’s Disease, Ataxia, Corticobasal Syndrome, Corticobasal Degeneration, Cervical Dystonia, Gerstmann Schenker Schyuler Syndrome, Wilson Disease, MSA-Parkinsonian, and Essential Tremor. In this study, this group is considered an active control group against the PD group.
@@
Category Number of % Females Age Files MoCA UPDRS3
@@ II-B1 Point Tasks
[…]
@@ II-B3 Writing Tasks
Participants first engaged in the Copy Text task, wherein they were tasked with replicating a paragraph of text displayed on the screen. We then focus on the impact of multitasking by introducing the Copy Read Text task, where participants copied another text while simultaneously reading it aloud. Subsequently, participants’ spontaneous writing abilities were captured through the Freewrite task, prompting them to write freely one or two sentences that must not contain any personal information. Lastly, we exploring handwriting a different context, the Numbers task, wherein participants were tasked with solving eight simple arithmetic operations summing two natural numbers not exceeding a total of 20, while saying the results of the sums out loud. It is worth noting that the Numbers task may be analyzed differently from the writing tasks due to its repetitive nature, necessitating a distinct analytical approach.
@@ III-B Task-Agnostic Metrics
    segmentation of each series using only the segments gathered In-Air or On-Tablet, and additionally decomposing the On-Tablet sequence in a suite of strokes. The segmentation is described in Section III-B2
@@
Figure 3: Pipeline used to extract task-agnostic metrics across all tasks.
@@ III-B3 Time-Independent Metrics
We chose not to utilize the amplitude measure for angular velocity and accelerations, as it would always yield identical values. Furthermore, we calculated the Shannon Entropy based on the points’ positions, whether on the tablet or in the air. To achieve this, we employ a Gaussian kernel density estimate (KDE) on the points’ $X$ and $Y$ positions. This enabled us to determine the probability $q_{i}$ for each point $(x_{i},y_{i})$. Shannon Entropy is then defined as $H_{s}=-\sum_{i=1}^{N}q_{i}ln(q_{i})$. Following the computation of these metrics, we obtained a total of 54 task-agnostic metrics across the 14 tasks.
@@ III-C Task-Specific Metrics
In this section, we detail the various task-specific metrics we computed. We divided the presentations of the metrics along the four main tasks’ categories proposed in Section II-B, with the addition of a section specific to the Numbers task.
@@ III-C1 Point Tasks Metrics
This section presents the metrics specific to the Point Tasks. Figure 4 depicts these metrics. Point Tasks are mainly used to measure the motor stability of participants and are mostly used for characterization of tremor, primarily relevant to PD or PDM (though tremor may also be observed in AD, due to copathology). Most dynamic metrics have already been explored as task-agnostic metrics, so we focused on the static metrics of the drawing: how symmetric and regular it is, and what the dispersion of the samples is. For this task, we calculated the geometric center of the pen’s position over the digital tablet. We used this geometric center as a pivot to compare the distribution of the points with respect to the computed center, through 3 metrics:
@@
    Average Angle measures the average of the angle between a point and the center, which can be used as a proxy for the symmetry of the drawing. Tremor and bradykinesia in PD do not always affect all limb regions equally. An average angle ¿ 0 might implicate a strong tremor axis.
@@ III-C2 Spiral Tasks Metrics
  • Eccentricity: $Ec=\sqrt{1+\frac{b^{2}}{a^{2}}}$. The eccentricity is a scalar [0,1[$\in$[0,1[∈ [ 0 , 1 [ that measures the compression of an ellipse when defined as a conic section. An eccentricity of 0 would be the one of a circle, while an eccentricity of 1 would indicate a parabola.
@@ III-C3 Numbers Task Metrics
The average answer pressure metric is a measure of the amount of pressure that a person applies while writing the answer. This metric can be leveraged as a time series metric to track its variation as the participant progresses through the task. It is computed through a linear regression between the average pressure of each answer and the question number. The metric itself represents the slope of this regression line. This metric might help pick up dystonia features, involuntary muscle contractions commonly seen in neurodegenerative diseases.
@@
The total duration metric measures the time a participant spends answering one of the questions. Given the absence of a predefined time limit for the Numbers task, this metric provides insight into the participant’s pace in completing the task.
@@ III-C4 Writing Tasks Metrics
Figure 7: Schematic of the metrics used in writing tasks, illustrated for the words ’This is an example’.
@@
The total time metric gauges the duration required for task completion. It is determined by calculating the time elapsed from when the participant initially places the pen to start writing until they lift the pen after writing the last word. This metric holds greater significance for tasks like Copy Text and Copy Read Text, where participants’ required text to write is the same.
@@ III-D Statistical Analysis
After extracting a total of 76 metrics (i.e., 54 task-agnostic metrics and 22 task-specific metrics), we measure which can be used to assess and characterize NDs. Our goal is twofold: firstly, to discern which metrics can effectively differentiate between the CTL group and the other experimental groups under consideration, and secondly, to ascertain whether handwriting metrics alone can distinguish PD from other forms of Parkinsonism (PDM). To achieve this, we conduct Kruskal’s H-test across all possible pairs of groups for each computed metric, using the Scipy toolkit [^2] https://docs.scipy.org/doc/scipy/reference/generated/scipy.stats.kruskal.html. Specifically, we examine the ability of each metric to distinguish the CTL group from each disease group, which requires the following comparisons: AD vs. CTL, PD vs. CTL, and PDM vs. CTL. We also assess the effectiveness of each metric in differentiating the PD from the PDM group. This procedure yields four comparisons, resulting in four p-values per metric.
The underlying experimental hypothesis is that the information in a given task can be leveraged to extract descriptors that can discriminate between two groups. To avoid false discovery, we apply FDR correction (from the Statsmodel toolkit [^3] https://www.statsmodels.org/dev/generated/statsmodels.stats.multitest.fdrcorrection.html on all the p-values grouped by task and test, so we apply $4\times 14=56$ corrections. Significant results from the statistical analysis are presented and commented on in the following section.
To complete our analysis of the separability of those groups of participants, we measured the Area Under the Curve (AUC) for pairs of participant groups, which will be presented alongside FDR-corrected p-values.
III-E Additional comparisons
The analyzed tasks consist of a carefully selected subset of related activities, allowing for a comparative analysis of a specific task group. Two types of comparisons are proposed:
  • Grouped metrics: Instead of measuring the significance of the metrics individually, we compare metrics for groups of tasks, such as the average speed, for all spirals.
  • Differences of metrics: Instead of considering the metrics, we consider the difference between the same metric evaluated on 2 different tasks.
Figure 8 shows how we perform those operations for the different tasks.
Figure 8: Schematic of the organization of the tasks when comparing 2 tasks or grouping a set of tasks.
III-E1 Grouped metrics

### v3

@@ Abstract
Motor dysfunction is a common sign of neurodegenerative diseases (NDs) such as Parkinson's disease (PD) and Alzheimer’s disease (AD), but may be difficult to detect, especially in the early stages. In this work, we examine the behavior of a wide array of explainable metrics extracted from the handwriting signals of 113 subjects performing multiple tasks on a digital tablet, as part of the Neurological Signals dataset. The aim is to measure their effectiveness in characterizing NDs, including AD and PD. To this end, task-agnostic and task-specific metrics are extracted from 14 distinct tasks. Subsequently, through statistical analysis and a series of classification experiments, we investigate which metrics provide greater discriminative power between NDs and healthy controls and amongst different NDs. Preliminary results indicate that the tasks at hand can all be effectively leveraged to distinguish between the considered set of NDs, specifically by measuring the stability, the speed of writing, the time spent not writing, and the pressure variations between groups from our handcrafted explainable metrics, which shows $p$-values lower than 0.0001 for multiple tasks. Using various binary classification algorithms on the computed metrics, we obtain up to 87% accuracy for the discrimination between AD and healthy controls (CTL), and up to 69% for the discrimination between PD and CTL.
@@
Handwriting Analysis, Alzheimer's Disease, Parkinson's Disease, Digital Biomarkers, Neurological Signals
@@ I Introduction
Neurodegenerative diseases (NDs), including Parkinson’s disease (PD) and Alzheimer’s disease (AD), progressively impair cognitive, motor, and behavioral functions by affecting both overlapping and distinct brain regions. To monitor these dysfunctions, many studies have explored non-invasive techniques using wearable or digital sensors. Among them, handwriting stands out as a promising modality due to its intricate integration of cognitive, motor, and perceptual processes [1, 2]. Handwriting production recruits a distributed neural network across cortical and subcortical structures, including regions involved in learning, executive function, praxis, vigilance, and motor coordination—spanning the cerebral cortex, basal ganglia, and cerebellum [3]. Consequently, abnormalities in handwriting patterns can serve as sensitive digital biomarkers of disease presence and progression [4], with micrographia being a well-established early symptom of PD [5].
Both paper-based and digital handwriting tasks have been employed to evaluate ND-related impairments by analyzing the spatial and temporal characteristics of handwriting signals [4, 5, 1]. Historically, most studies have focused on PD, largely due to the overt motor symptoms associated with the disease, including micrographia [6, 7, 8, 9], tremor [10], and bradykinesia [10]—collectively referred to as PD dysgraphia [1]. Handwriting analysis has also been explored for AD detection [11], as the disease’s prevalence continues to rise [12, 13, 14].
Numerous handwriting datasets have been introduced to study NDs, capturing either static images or dynamic time-series signals. However, most datasets are restricted to a single disease type and a single data modality, thereby limiting their experimental scope. For PD: HandPD [15]: Includes 92 participants (74 PD, 18 CTL), with image-based recordings of spiral and meander tracing. Naive Bayes classification achieved up to 78.9% accuracy. NewHandPD [15]: Expanded the previous dataset by adding pressure, acceleration, and tilt data from 14 PD and 21 CTL participants. Using CNNs, the study reached 83.77% accuracy on spirals and 87.14% on meanders. PaHaW [6]: Collected position and pressure data from spirals and repeated text in 37 PD and 38 CTL subjects. Using explainable metrics and SVMs, it achieved 81.3% accuracy. ParkinsonHW [16]: Recorded dynamic handwriting from 62 PD and 15 CTL subjects via a digital tablet. Cascaded non-neural classifiers achieved 84.67% accuracy on spirals and 90.91% on stability tasks [17].
For AD: Early work by [14] evaluated handwriting signatures from 29 AD and 30 CTL subjects, achieving a 3% error rate with a decision tree model. The dataset introduced by [13] included 23 AD, 12 MCI, and 17 CTL participants performing drawing and writing tasks on a tablet. Sentence copying reached up to 80.4% classification accuracy across three groups. The DARWIN dataset [12], now a standard benchmark, includes 100 AD and 90 CTL participants engaged in copying, memory, and graphical tasks. Fusing non-neural classifiers trained on individual tasks yielded a top accuracy of 94.29%.
Handwriting tasks for ND assessment can be categorized by complexity and task structure: Simple motor tasks involve drawing basic shapes such as spirals or straight lines, or stability exercises, such as maintaining a fixed pen over a tablet. These evaluate tremor, trajectory smoothness, velocity, and acceleration [18]. Written language tasks involve producing words or sentences. On tablets, these tasks enable the study of in-air movement and motor planning between words [18]. In-air time reflects cognitive preparation [5], and since in-air and on-surface dynamics are statistically independent [19], both provide complementary information. Increased in-air duration may reflect hesitation or planning deficits. Complex visuoconstructive tasks include copying figures or drawing clocks. The clock drawing test is an established tool for evaluating visuospatial and executive function across several NDs [20, 21]. Copying tasks allow examination of written response fidelity in varied contexts [22].
In this work, we introduce a new dataset composed of digital handwriting samples from 103 participants, including individuals diagnosed with AD, Mild Cognitive Impairment (MCI), PD, and Related Disorders (a heterogeneous group we refer to as PDM, comprising atypical Parkinsonism and misdiagnosed cases). Participants completed 14 tablet-based handwriting tasks specifically designed to probe cognitive and motor deficits across these conditions. We extract a large set of explainable, quantitative metrics, referred to as handwriting metrics, from these tasks. These include both: Task-agnostic metrics, which capture global signal properties across all tasks, regardless of their nature, and Task-specific metrics, which target metrics tailored to particular task types. We then assess which metrics best distinguish between groups through statistical analyses and binary classification experiments. In particular, we compare ND subgroups (PD, AD, PDM) to healthy controls, and contrast PD and PDM to identify discriminative metrics within motor phenotypes.
The main contributions of this paper are as follows:
  • Introduction of a novel, multi-condition handwriting dataset for ND assessment.
  • Inclusion of multiple disease groups, including an active control group (PDM), expanding beyond the narrower focus of prior datasets.
  • Extension of handwriting analysis from PD to AD and MCI, which have received limited attention.
  • Design and evaluation of a broad set of task-agnostic and task-specific metrics that align with expected behavioral patterns and reveal significant group differences.
  • Novel exploration of inter-task comparisons to analyze behavioral shifts when task parameters change (e.g., drawing hand).
@@ II-A Data Collection
The authors of this study collected a dataset from 113 participants from a larger digital biomarker cohort, NeuroLogical Signals, which our group has previously reported [23, 24, 25, 26, 27, 28, 29]. These participants were either categorized as healthy (CTL) or diagnosed and treated for a ND by a subspecialist neurologist or geriatric psychiatrist from Johns Hopkins Hospital. Participants diagnosed with NDs were categorized into aforementioned groups based on disease specific clinical diagnostic criteria diagnosis:
@@
  • PD participants met international recognized diagnostic criteria at a "clinically established" [33] level of certainty and were followed by a movement disorders neurologist.
  • PDM participants presented with one or more core symptoms of PD, meeting a "Probable PD" diagnosis but which was later revised based on evolving characteristics or supportive testing. This is a mixed, active comparator group includes patients later diagnosed with Dementia with Lewy Bodies, Fahr’s Disease, Ataxia, Corticobasal Syndrome, Corticobasal Degeneration, Cervical Dystonia, Gerstmann Schenker Schyuler Syndrome, Wilson Disease, MSA-Parkinsonian, and Essential Tremor. In this study, this group is considered an active control group against the PD group.
@@
Category Number of % Females Age Files MoCA [34] UPDRS3 [35]
@@ II-B1 Point Tasks
The point tasks are designed to assess the static motor control and stability of the participants. Each participant was asked to maintain the pen vertically above a point shown in the tablet without moving or touching the tablet for $10\text{\,}\sec$. More specifically, in the Point Dominant and Point Non-Dominant tasks, participants are asked to use their dominant and Non-Dominant hand respectively, and in the Point Sustained, they are asked to use their dominant hand, in addition, to sustain the phonation of the vowel A during the exercise.
@@ II-B3 Writing Tasks
Participants first engaged in the Copy Text task, wherein they were tasked with replicating a paragraph of text displayed on the screen. We then focus on the impact of multitasking by introducing the Copy Read Text task, where participants copied another text while simultaneously reading it aloud. Subsequently, participants' spontaneous writing abilities were captured through the Freewrite task, prompting them to write freely one or two sentences that must not contain any personal information. Lastly, we exploring handwriting a different context, the Numbers task, wherein participants were tasked with solving eight simple arithmetic operations summing two natural numbers not exceeding a total of 20, while saying the results of the sums out loud. It is worth noting that the Numbers task may be analyzed differently from the writing tasks due to its repetitive nature, necessitating a distinct analytical approach.
@@ III-B Task-Agnostic Metrics
    segmentation of each series using only the segments gathered In-Air or On-Tablet, and additionally decomposing the On-Tablet sequence in a suite of strokes. A stroke being defined as the trajectory On-Tablet between two In-Air moments. The segmentation is described in Section III-B2
@@
Figure 3: Pipeline used to extract task-agnostic metrics across all tasks. The amplitudes for angular speech and acceleration are not kept, which is why we have 54 and not 58 metrics.
@@ III-B3 Time-Independent Metrics
We chose not to utilize the amplitude measure for angular velocity and accelerations, as it would always yield identical values. Furthermore, we calculated the Shannon Entropy based on the points' positions, whether on the tablet or in the air. To achieve this, we employ a Gaussian kernel density estimate (KDE) on the points' $X$ and $Y$ positions. This enabled us to determine the probability $q_{i}$ for each point $(x_{i},y_{i})$. Shannon Entropy is then defined as $H_{s}=-\sum_{i=1}^{N}q_{i}ln(q_{i})$. Following the computation of these metrics, we obtained a total of 54 task-agnostic metrics across the 14 tasks.
@@ III-C Task-Specific Metrics
In this section, we detail the various task-specific metrics we computed. We divided the presentations of the metrics along the four main tasks' categories proposed in Section II-B, with the addition of a section specific to the Numbers task.
@@ III-C1 Point Tasks Metrics
This section presents the metrics specific to the Point Tasks. Figure 4 depicts these metrics. Point Tasks are mainly used to measure the motor stability of participants and are mostly used for characterization of tremor, primarily relevant to PD or PDM (though tremor may also be observed in AD, due to copathology). Most dynamic metrics have already been explored as task-agnostic metrics, so we focused on the static metrics of the drawing: how symmetric and regular it is, and what the dispersion of the samples is. For this task, we calculated the geometric center of the pen's position over the digital tablet. We used this geometric center as a pivot to compare the distribution of the points with respect to the computed center, through 3 metrics:
@@
    Average Angle measures the average of the angle between a point and the center, which can be used as a proxy for the symmetry of the drawing. Tremor and bradykinesia in PD do not always affect all limb regions equally. An average angle >>>> 0 or <<0<<0 might implicate a strong tremor axis.
@@ III-C2 Spiral Tasks Metrics
  • Eccentricity: $Ec=\sqrt{1+\frac{b^{2}}{a^{2}}}$. The eccentricity is a scalar $\in[0,1[$ that measures the compression of an ellipse when defined as a conic section. An eccentricity of 0 would be the one of a circle, while an eccentricity of 1 would indicate a parabola.
@@ III-C3 Numbers Task Metrics
The average answer pressure metric is a measure of the amount of pressure that a person applies while writing the answer. This metric can be leveraged as a time series metric to track its variation as the participant progresses through the task. It is computed through a linear regression between the average pressure of each answer and the question number. The metric itself represents the slope of this regression line. This metric might help pick up dystonia metrics, involuntary muscle contractions commonly seen in neurodegenerative diseases.
@@
The total duration metric measures the time a participant spends answering one of the questions. Given the absence of a predefined time limit for the Numbers task, this metric provides insight into the participant's pace in completing the task.
@@ III-C4 Writing Tasks Metrics
Figure 7: Schematic of the metrics used in writing tasks, illustrated for the words 'This is an example'.
@@
The total time metric gauges the duration required for task completion. It is determined by calculating the time elapsed from when the participant initially places the pen to start writing until they lift the pen after writing the last word. This metric holds greater significance for tasks like Copy Text and Copy Read Text, where participants' required text to write is the same.
@@ III-D Statistical Analysis
After extracting a total of 76 metrics (i.e., 54 task-agnostic metrics and 22 task-specific metrics), we aim to evaluate their relevance for assessing and characterizing neurodegenerative diseases (NDs). Our objective is twofold: first, to identify which metrics effectively differentiate the CTL group from each of the disease groups (AD, PD, and PDM), and second, to assess whether handwriting metrics can distinguish PD from other forms of Parkinsonism (PDM).
To this end, we perform a Kruskal-Wallis H-test across the four experimental groups (CTL, AD, PD, PDM) for each metric, using the scipy.stats.kruskal function from the Scipy toolkit [^2] https://docs.scipy.org/doc/scipy/reference/generated/scipy.stats.kruskal.html. For metrics showing a significant global effect (p<<0.05), we apply Dunn’s post hoc test to identify specific group differences. To control for false discoveries arising from multiple testing, we apply False Discovery Rate (FDR) correction using the statsmodels.stats.multitest.fdrcorrection function [^3] https://www.statsmodels.org/dev/generated/statsmodels.stats.multitest.fdrcorrection.html across the set of p-values grouped by task and test. Since we test 14 tasks across 4 pairwise group comparisons, we apply $4\times 14=56$ FDR corrections. To complete our analysis of the separability of those groups of participants, we measured the Area Under the Curve (AUC) for pairs of participant groups, which will be presented alongside FDR-corrected p-values.
Additionally, to investigate how individual metrics relate to cognitive and motor functioning, we compute the Pearson correlation coefficient between each of the 76 extracted features and both the MoCA [34] (Montreal Cognitive Assessment, for all participants) and UPDRS-III [35] (Unified Parkinson's Disease Rating Scale, motor section, only measured for the PD and PDM participants) scores across participants. To provide a more detailed analysis of motor function, we also compute correlations between each feature and the individual subdomains of the UPDRS-III score, including: rigidity, upper extremity bradykinesia, lower extremity bradykinesia, arising from a chair, gait, freezing of gait, postural stability, posture, gait and posture composite, kinetic tremor, postural tremor, resting tremor, global tremor score, and non-tremor components.
Significant results from both the group comparisons and correlation analyses are presented and discussed in the following sections.
III-D1 Grouped metrics